OpenStudy (anonymous):

Simplify: (i) (a^3 b^2 c^4)/(ab^3 c^(-3) )

myininaya (myininaya):

You would like to simplify I bet :) Now you need to recall the law of exponents! \[\frac{x^a}{x^b}=x^{a-b}\] ; \[ x^{a} x^{b}=x^{a+b} \]

myininaya (myininaya):

\[\frac{a^3 b^2 c^4}{ab^3c^{-3}}=\frac{a^3}{a} \cdot \frac{b^2}{b^3} \cdot \frac{c^4}{c^{-3}}\]

myininaya (myininaya):

So using just the first thing I wrote rewrite this

myininaya (myininaya):

How about I rewrite just a little so you can try to rewrite my writing of it \[\frac{a^3}{a^1} \cdot \frac{b^2}{b^3} \cdot \frac{c^4}{c^{-3}}\] So for that first little part \[\frac{a^3}{a^1}=a^{3-1}=a^2\] What do you think if we apply that first rule here: \[\frac{b^2}{b^3}\]